When 4.00 g of sodium hydrogen carbonate reacts completely with 90.0 mL of

acetic acid, the temperature increases from 20.00 ºC to 40.00 ºC. The products are

carbon dioxide, water and sodium acetate. Calculate the molar enthalpy of reaction,

ΔHrxn. Include a balanced thermochemical reaction.


1
Expert's answer
2021-02-11T04:52:58-0500

Let us assume that the dencity and the heat capacity of the solution are equal to the dencity and the heat capacity of water and does not depend on the temperature:

Csol = 4.18 (kJ/(kg × K))

msol = 0.090 + 0.004 = 0.094 (kg)

"\Delta"Q = 4.18 × 0.094 × (40-20) = 7.85 (kJ)

n = m / M

n(NaHCO3) = 4 / 84 = 0.0476 (mol) = n(H2O)

"\Delta"Hreaction = -7.85 / 0.0476 = -164.9 (kJ / mol)

NaHCO3 + CH3COOH = CH3COONa + H2O + CO2 + 164.9 kJ